OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = 2x2 – 8x + 6?

OpenStudy (anonymous):

For the y-intercept, plug in 0 for x. For the x-intercepts, plug in zero for y and factor it.

OpenStudy (amorfide):

factorise to get x factorising gives 2x2 -6x-2x+6 2x(x-3)-2(x-3) (2x-2)(x-3) x=2/3 or x=3 sub both into the original equation to get 2 y values this gives two coordinates

OpenStudy (anonymous):

0 = 2(x^2 - 4x + 3) 0 = (x - 3)(x - 1) x = 3 and 1.

OpenStudy (amorfide):

yeah you are right i got 2/2 not 2/3 lol

OpenStudy (anonymous):

Therefore, the y-intercept is 6. (0, 6) The x-intercepts are 3 and 1. (3, 0) (1, 0)

OpenStudy (anonymous):

for the answer choices there is no y-intercept that =-6

OpenStudy (anonymous):

I changes it

OpenStudy (anonymous):

i see now lol.

OpenStudy (anonymous):

can you help me on another one?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Describe the differences between the graph of y= -3/4(x-6)^2+4 and the standard position graph of y = x^2.

OpenStudy (anonymous):

Wait, are you in algebra I or algebra II? THat way I know how in depth I should go.

OpenStudy (anonymous):

alg. 2